# A Fast and Lightweight Model for Causal Audio-Visual Speech Separation

### v1

@@ 1 Introduction
With the rapid advancement of speech processing and artificial intelligence technologies, audio-only speech separation (AOSS) methods are facing numerous challenges in the cocktail party scenario [26]. These challenges are mainly manifested as a significant decline in processing efficiency with an increasing number of speakers, and insufficient robustness in complex noisy environments [29, 8]. Inspired by the human multisensory information processing mechanism [25], researchers have proposed strategies for integrating visual information with auditory information. By incorporating visual cues into the separation network, the robustness and processing efficiency of speech separation methods can be significantly improved [19, 18]; Such methods are referred to as audio-visual speech separation (AVSS).
However, most existing AVSS methods are non-causal and require the entire audio sequence as input, and their complex architectures incur large Params and MACs, limiting them to offline scenarios due to their inability to process streaming data in real time [21]. Therefore, developing low-latency causal models for real-time applications is highly demanded.
@@
Extensive experiments on mainstream datasets LRS2 [1], LRS3 [2], and VoxCeleb2 [6] demonstrated that Swift-Net achieved SOTA performance. Additionally, we provided a toolkit that modularizes the conversion of mainstream non-causal AVSS methods into causal methods. Source code was made available at https://anonymous.4open.science/r/Swift-Net.
@@ 2.1 Non-causal AVSS Model
In recent years, most AVSS models with decent separation performance have adopted non-causal architectures [34, 18, 19, 1, 9, 20, 30, 27, 28], achieving notable improvements in separation performance by integrating both historical and future information. Researchers have developed a variety of model architectures tackling the AVSS task, including convolutional neural networks (CNNs), recurrent neural networks (RNNs), Transformers, and hybrid ones. Specifically, CNN-based models (e.g., AV-ConvTasNet [34] and CTCNet [19]) typically utilize time-domain convolutional networks and incorporate visual lip-reading features to enhance separation performance. However, such methods primarily perform audio-visual fusion through convolutional operations, resulting in large model parameter count and high computational complexity. Furthermore, they rely on future time-step information, making them challenging to deploy in real-time processing scenarios. RNN-based models (e.g., AV-DPRNN [30]) leverage deep recurrent structures to capture long-term dependencies. However, due to their sequential processing mechanism [15], their computational efficiency is limited. Transformer-based models, represented by AV-SepFormer [20], achieve synchronous modeling of modal information through a dual-path Transformer architecture, which substantially improves separation performance. Nonetheless, the computational complexity of such models scales quadratically along with sequence length, leading to increased computational burden and model size. Lastly, hybrid architecture models (e.g., AV-TFGridNet [27] and RTFS-Net [28]) integrate CNN, Transformer, and RNN modules to balance between performance and computational complexity, achieving superior separation results in offline speech separation tasks. However, these methods typically rely on global information, thus limiting their application in real-time separation scenarios. In this paper, we aim to develop a causal AVSS method capable of operating under streaming processing conditions.
@@ 3 Causal Design Strategies
Causal Average Pooling Layer. We propose a segment-based causal adaptive average pooling method, as shown in Figure 1. Specifically, let the input sequence be $\mathbf{X}\in\mathbb{R}^{1\times T}$, where $T$ denotes the total number of timesteps. We first uniformly divide $\mathbf{X}$ along the temporal axis into $N$ non-overlapping segments, each with length $L=\lceil T/N\rceil$. For the $n$-th output interval ($n\in\{1,2,\dots,N\}$), the corresponding temporal range is $[(n-1)L+1,\,nL]$. To satisfy the causality constraint during the pooling operation, the pooling result of the $n$-th segment, $\mathbf{y}_{n}$, depends only on the current and previous inputs, that is,
@@ 4 Method
Our proposed causal AVSS model, Swift-Net, consists of an audio encoder, a video encoder, a separation module, and an audio decoder, as illustrated in Figure 2. Specifically, the input comprises a mixed audio signal $\mathbf{Y}\in\mathbb{R}^{1\times T_{a}}$ and the corresponding grayscale lip movement video sequence $\mathbf{V}\in\mathbb{R}^{1\times T_{v}\times H\times W}$, where $T_{a}$ denotes the length of the audio sequence, and $T_{v}$, $H$, and $W$ represent the number, height, and width of video frames, respectively. First, the audio encoder maps the mixed audio signal to an audio embedding $\mathbf{E_{a}(0)}\in\mathbb{R}^{C_{a}\times T^{\prime}_{a}\times F}$, and the video encoder maps the lip movement video sequence to a lip embedding $\mathbf{E}_{v}\in\mathbb{R}^{C_{v}\times T_{v}}$, where $C_{a}$ and $C_{v}$ denote the dimensions of the encoded features, and $T^{\prime}_{a}$ is the corresponding encoded temporal lengths, $F$ denotes the number of frequency bins in the spectrogram. Subsequently, a causal audio-visual separation network generates an estimated mask $\mathbf{M}\in\mathbb{R}^{C_{a}\times T^{\prime}_{a}\times F}$ for the target speaker based on $\mathbf{E_{a}(0)}$ and $\mathbf{E}_{v}$. Finally, following the approach of RTFSNet [28], the audio decoder performs element-wise multiplication of $\mathbf{E_{a}(0)}$ and $\mathbf{M}$ in the complex domain to obtain the audio embedding of the target speaker, $\mathbf{R}_{a}\in\mathbb{R}^{C_{a}\times T^{\prime}_{a}\times F}$. Finally, $\mathbf{R}_{a}$ is transformed back to the time-domain waveform by applying the decoder, thus yielding the separated speech of the target speaker $\bar{\mathbf{A}}\in\mathbb{R}^{1\times T_{a}}$.
@@ 4.1 Encoder
where $F$ and $T^{\prime}_{a}$ denote the frequency and temporal dimensions, respectively. To further enrich the audio representation, $\mathbf{G}$, $\mathbf{Y}_{r}$, and $\mathbf{Y}_{i}$ are stacked along the channel dimension to construct a composite audio feature $\mathbf{Y}_{m}\in\mathbb{R}^{3\times T^{\prime}_{a}\times F}=\{\mathbf{G}||\mathbf{Y}_{r}||\mathbf{Y}_{i}\}$, where ||||| | denotes concatenation along the channel dimension. Thereafter, $\mathbf{Y}_{m}$ is fed into a convolutional layer with normalization and nonlinear activation to obtain the audio embedding $\mathbf{{E_{a}}(0)}$, which is subsequently input to the following audio-visual separation network.
@@ 4.2 The Separation Module
The overall framework comprises three core components: a lightweight video processing block (LightVid block), an efficient audio block (FTGS block), and a selective attention audio-visual fusion block (SAF block). First, visual features $\mathbf{E}_{v}$ generated by the video encoder are fed into the LightVid block to obtain enhanced visual embeddings $\bar{\mathbf{E}}_{v}$. Simultaneously, audio features $\mathbf{E_{a}(0)}$ are processed by the FTGS block to generate refined audio embeddings $\mathbf{E_{a}(1)}$. Subsequently, features from both modalities $\{\bar{\mathbf{E}}_{v}\in\mathbb{R}^{C_{v}\times T_{v}},\mathbf{{E}_{a}(1)}\in\mathbb{R}^{C_{a}\times T^{\prime}_{a}\times F}\}$ are input into the SAF block to produce a joint audio-visual representation $\mathbf{{E}_{av}}\in\mathbb{R}^{C_{a}\times T^{\prime}_{a}\times F}$. Next, $\mathbf{\mathbf{E}_{av}}$ is fed into a deep network composed of $N-1$ stacked Efficient FTGS blocks, where each FTGS block shares parameters to further integrate and refine the joint representation, ultimately yielding iteratively optimized features $\mathbf{{E}_{a}(N)}\in\mathbb{R}^{C_{a}\times T^{\prime}_{a}\times F}$. This cascaded structure, combined with the parameter sharing mechanism, significantly enhances the computational efficiency and effectively reduces the model size, enabling the separator to achieve real-time, low-latency inference while maintaining superior separation performance.
@@ 4.2.1 The LightVid block
Most of the current mainstream AVSS models employ complex visual encoders involving delicate usage of deep convolutional neural networks or Transformer [31] architectures, often entailing considerable computational cost. To address the this issue, we propose a lightweight video processing module, termed LightVid block (see Figure 3), which serves as an efficient causal visual feature extractor capable of extracting both local spatial information and long-range temporal dynamics from lip-reading video sequences with low latency.
@@ 4.2.2 The FTGS block
Subsequently, for $\mathbf{A}_{f}$, unfolds along the time axis, and the above procedure is repeated using Bidirectional Grouped SRUs (introduced later) to capture the temporal dynamics of the feature sequences. In contrast to frequency modeling, temporal modeling employs unidirectional grouped SRUs to ensure a causal structure, enhancing the integration of information from historical feature spaces and thereby improving separation performance. The output is denoted as $\mathbf{A}_{t}$. Next, a causal self-attention mechanism is applied to $\mathbf{A}_{t}$ to further improve the audio feature representation, resulting in $\mathbf{A}_{att}$. Finally, a reconstruction strategy similar to TDANet [17] is adopted to recover the time-frequency resolution, generating the next-level feature $\mathbf{E}_{a}(i+1)$.
@@ 5.2 Implementation Details
For training, we employed an early stopping strategy with a maximum number of 500 epochs. All experiments were conducted on 8 NVIDIA H800 GPUs with batch size 8. The optimizer used was AdamW [22] with weight decay $1\times 10^{-1}$ and initial learning rate $1\times 10^{-3}$. If the validation loss did not reach new low for 5 consecutive epochs, the learning rate was halved. The loss function was defined as the scale-invariant signal-to-noise ratio (SI-SNR) [14] between the predicted and target speech signals.
@@
5.3 Comparison with Existing Methods and Causal AVSS Models
We conducted a systematic evaluation of Swift-Net on three datasets and compared it with various mainstream AVSS models. Results are shown in Table 1. To ensure fairness, all comparison methods were adapted to their causal versions following the causal strategy described in Section 3, and the open-source codes of each model were available [^2] https://anonymous.4open.science/r/Swift-Net. Moreover, all compared models were trained within the same training framework and evaluated using consistent performance metrics. Across all datasets, it was noteworthy that Swift-Net-6, as the fastest inference variant, achieved an SI-SNRi of 13.4 dB, outperforming existing SOTA causal methods such as causal AV-TF-GridNet. In addition, Swift-Net-6 reduced the parameter count by 11.8 times and the computational cost by 10.3 times compared with causal AV-TF-GridNet. Furthermore, our high-performance variant, Swift-Net-12, achieved over 1 dB improvement in performance relative to causal AV-TF-GridNet. These results collectively demonstrated that Swift-Net offered significant advantages for causal speech separation tasks.
In addition, we further conducted a comparative analysis of the separation performance of RTFSNet [28], AV-TF-GridNet [27], and Swift-Net in real-world multi-speaker scenarios. Specifically, we collected four distinct multi-speaker video samples from the YouTube platform. To facilitate intuitive comparison of the separated speech results, we developed an interactive demo website [^3] https://anonymous.4open.science/w/Swift-Net-Demo/ allowing users to listen and compare the outputs. Subjective listening evaluations indicated that, compared to RTFSNet and AV-TF-GridNet, Swift-Net produces speech signals that were clearer, more natural, effectively suppressing background interference.
@@ 5.4 Ablation Studies
Different Visual Feature Extraction Modules. We replaced the LightVid block visual-feature encoder with those from ConvTasNet and RTFSNet while keeping every other component of the streaming AVSS pipeline identical in order to quantify its module-level benefits. As shown in Table 2, this lean design requires only 67.27 K parameters and 3.39 M MACs – over 4× reduction relative to ConvTasNet and RTFSNet. Remarkably, despite its minimal complexity, LightVidBlock still delivers a slight performance edge, confirming that a structurally simple encoder can unite efficiency and effectiveness in real-time causal AVSS.
@@
Effect of Incorporating Power Spectrogram Features. We systematically evaluated the impact of introducing audio power spectrogram features on audio separation performance across three models: RTFSNet, AV-TF-GridNet, and Swift-Net. As shown in Table 6, in Swift-Net (with 0.53M parameters and an increase of less than 0.1G MACs), the inclusion of power spectrogram features raised the SI-SNRi from 13.1 to 13.3. Furthermore, introducing power spectrogram features resulted in negligible increases in model parameters and incurred minimal MACs overhead, making it a low-cost enhancement. We conjectured power spectrogram features provided key speech-presence cues, thereby improving audio-visual fusion and separation performance.
Table 6: Results of with (upper) and without (lower) power spectrogram modeling on the LRS2-2Mix dataset.
@@
RTFSNet 12.8 12.4 13.1 12.8 0.64 0.63 49.89 49.81
AV-TF-GridNet 13.0 12.7 13.4 13.1 6.26 6.25 213.96 213.88
Swift-Net-6 13.3 13.1 13.6 13.4 0.53 0.52 20.68 20.60

### v2

@@ 1 Introduction
With the rapid advancement of speech processing and artificial intelligence technologies, audio-only speech separation (AOSS) methods are facing numerous challenges in the cocktail party scenario [29]. These challenges are mainly manifested as a significant decline in processing efficiency with an increasing number of speakers, and insufficient robustness in complex noisy environments [32, 10]. Inspired by the human multisensory information processing mechanism [28], researchers have proposed strategies for integrating visual information with auditory information. By incorporating visual cues into the separation network, the robustness and processing efficiency of speech separation methods can be improved [22, 21]; Such methods are referred to as audio-visual speech separation (AVSS).
However, most existing AVSS methods are non-causal and require the entire audio sequence as input, and their complex architectures incur large Params and MACs, limiting them to offline scenarios due to their inability to process streaming data in real-time [24]. Therefore, developing low-latency causal models for real-time applications is highly demanded.
@@
Extensive experiments on mainstream datasets LRS2 [1], LRS3 [2], and VoxCeleb2 [7] demonstrated that Swift-Net achieved SOTA performance. Additionally, we provided a toolkit that modularizes the conversion of mainstream non-causal AVSS methods into causal methods. Source code was made available at https://github.com/JusperLee/Swift-Net.
@@ 2.1 Non-causal AVSS Model
Early research on speech separation and enhancement largely centered on audio-only processing [36, 8, 26, 3, 15]. Many of these methods require pre-registered reference utterances to extract the target speaker, which limits their applicability in open-world scenarios. To mitigate these limitations, subsequent studies fused audio with visual cues, giving rise to audio-visual speech separation (AVSS). However, most AVSS models with decent separation performance have adopted non-causal architectures [38, 21, 22, 1, 11, 23, 33, 30, 31], achieving notable improvements in separation performance by integrating both historical and future information. Researchers have developed a variety of model architectures tackling the AVSS task, including convolutional neural networks (CNNs), recurrent neural networks (RNNs), Transformers, and hybrid ones. Specifically, CNN-based models (e.g., AV-ConvTasNet [38] and CTCNet [22]) typically utilize time-domain convolutional networks and incorporate visual lip-reading features to enhance separation performance. However, such methods primarily perform audio-visual fusion through convolutional operations, resulting in large model parameter count and high computational complexity. Furthermore, they rely on future time-step information, making them challenging to deploy in real-time processing scenarios. RNN-based models (e.g., AV-DPRNN [33]) leverage deep recurrent structures to capture long-term dependencies. However, due to their sequential processing mechanism [18], their computational efficiency is limited. Transformer-based models, represented by AV-Sepformer [23], achieve synchronous modeling of modal information through a dual-path Transformer architecture, which substantially improves separation performance. Nonetheless, the computational complexity of such models scales quadratically along with sequence length, leading to increased computational burden and model size. Lastly, hybrid architecture models (e.g., AV-TF-GridNet [30] and RTFS-Net [31]) integrate CNN, Transformer, and RNN modules to balance between performance and computational complexity, achieving superior separation results in offline speech separation tasks. However, these methods typically rely on global information, thus limiting their application in real-time separation scenarios. In this paper, we aim to develop a causal AVSS method capable of operating under streaming processing conditions.
@@ 3 Causal Design Strategies
Causal Average Pooling Layer. We propose a segment-based causal adaptive average pooling method, as shown in Figure 1. Specifically, let the input sequence be $\mathbf{X}\in\mathbb{R}^{1\times T}$, where $T$ denotes the total number of timesteps. We first uniformly divide $\mathbf{X}$ along the temporal axis into $N$ non-overlapping segments, each with length $L=\lceil T/N\rceil$. For the $n$-th output interval ($n\in\{1,2,\dots,N\}$), the corresponding temporal range is [(n1)L+1,nL][(n-1)L+1,\,nL]. To satisfy the causality constraint during the pooling operation, the pooling result of the $n$-th segment, $\mathbf{y}_{n}$, depends only on the current and previous inputs, that is,
@@ 4 Method
Our proposed causal AVSS model, Swift-Net, consists of an audio encoder, a video encoder, a separation module, and an audio decoder, as illustrated in Figure 2. Specifically, the input comprises a mixed audio signal $\mathbf{Y}\in\mathbb{R}^{1\times T_{a}}$ and the corresponding grayscale lip movement video sequence $\mathbf{V}\in\mathbb{R}^{1\times T_{v}\times H\times W}$, where $T_{a}$ denotes the length of the audio sequence, and $T_{v}$, $H$, and $W$ represent the number, height, and width of video frames, respectively. First, the audio encoder maps the mixed audio signal to an audio embedding $\mathbf{E_{a}(0)}\in\mathbb{R}^{C_{a}\times T^{\prime}_{a}\times F}$, and the video encoder maps the lip movement video sequence to a lip embedding $\mathbf{E}_{v}\in\mathbb{R}^{C_{v}\times T_{v}}$, where $C_{a}$ and $C_{v}$ denote the dimensions of the encoded features, and $T^{\prime}_{a}$ is the corresponding encoded temporal lengths, $F$ denotes the number of frequency bins in the spectrogram. Subsequently, a causal audio-visual separation network generates an estimated mask $\mathbf{M}\in\mathbb{R}^{C_{a}\times T^{\prime}_{a}\times F}$ for the target speaker based on $\mathbf{E_{a}(0)}$ and $\mathbf{E}_{v}$. Finally, following the approach of RTFSNet [31], the audio decoder performs element-wise multiplication of $\mathbf{E_{a}(0)}$ and $\mathbf{M}$ in the complex domain to obtain the audio embedding of the target speaker, $\mathbf{R}_{a}\in\mathbb{R}^{C_{a}\times T^{\prime}_{a}\times F}$. $\mathbf{R}_{a}$ is transformed back to the time-domain waveform by applying the decoder, thus yielding the separated speech of the target speaker $\bar{\mathbf{A}}\in\mathbb{R}^{1\times T_{a}}$.
@@ 4.1 Encoder
where $F$ and $T^{\prime}_{a}$ denote the frequency and temporal dimensions, respectively. To further enrich the audio representation, $\mathbf{G}$, $\mathbf{Y}_{r}$, and $\mathbf{Y}_{i}$ are stacked along the channel dimension to construct a composite audio feature $\mathbf{Y}_{m}\in\mathbb{R}^{3\times T^{\prime}_{a}\times F}=\{\mathbf{G}||\mathbf{Y}_{r}||\mathbf{Y}_{i}\}$, where $||$ denotes concatenation along the channel dimension. Thereafter, $\mathbf{Y}_{m}$ is fed into a convolutional layer with normalization and nonlinear activation to obtain the audio embedding $\mathbf{{E_{a}}(0)}$, which is subsequently input to the following audio-visual separation network.
@@ 4.2 The Separation Module
The overall framework comprises three core components: a lightweight video processing block (LightVid block), an efficient audio block (FTGS block), and a selective attention audio-visual fusion block (SAF block). First, visual features $\mathbf{E}_{v}$ generated by the video encoder feed into the LightVid block to obtain enhanced visual embeddings $\bar{\mathbf{E}}_{v}$. Simultaneously, audio features $\mathbf{E_{a}(0)}$ are processed by the FTGS block to generate refined audio embeddings $\mathbf{E_{a}(1)}$. Subsequently, features from both modalities $\{\bar{\mathbf{E}}_{v}\in\mathbb{R}^{C_{v}\times T_{v}},\mathbf{{E}_{a}(1)}\in\mathbb{R}^{C_{a}\times T^{\prime}_{a}\times F}\}$ flow into the SAF block to produce a joint audio-visual representation $\mathbf{{E}_{av}}\in\mathbb{R}^{C_{a}\times T^{\prime}_{a}\times F}$. Next, $\mathbf{\mathbf{E}_{av}}$ is fed into a deep network composed of N1N-1 stacked Efficient FTGS blocks, where each FTGS block shares parameters to further integrate and refine the joint representation, ultimately yielding iteratively optimized features $\mathbf{{E}_{a}(N)}\in\mathbb{R}^{C_{a}\times T^{\prime}_{a}\times F}$. This cascaded structure, combined with the parameter sharing mechanism, significantly enhances the computational efficiency and effectively reduces the model size, enabling the separator to achieve real-time, low-latency inference while maintaining superior separation performance.
@@ 4.2.1 The LightVid block
Most of the current mainstream AVSS models employ complex visual encoders involving delicate usage of deep convolutional neural networks or Transformer [34] architectures, often entailing considerable computational cost. To address this issue, we propose a lightweight video processing module, termed LightVid block (see Figure 3), which serves as an efficient causal visual feature extractor capable of extracting both local spatial information and long-range temporal dynamics from lip-reading video sequences with low latency.
@@ 4.2.2 The FTGS block
In contrast to frequency modeling, temporal modeling employs unidirectional grouped SRUs to ensure a causal structure, enhancing the integration of information from historical feature spaces and thereby improving separation performance. The output is denoted as $\mathbf{A}_{t}$. Next, a causal self-attention mechanism is applied to $\mathbf{A}_{t}$ to further improve the audio feature representation, resulting in $\mathbf{A}_{att}$. Finally, a reconstruction strategy similar to TDANet [20] is adopted to recover the time-frequency resolution, generating the next-level feature $\mathbf{E}_{a}(i+1)$.
@@ 5.2 Implementation Details
For training, we employed an early stopping strategy with a maximum number of 500 epochs. All experiments were conducted on 8 NVIDIA H800 GPUs with batch size 8. The optimizer used was AdamW [25] with weight decay $1\times 10^{-1}$ and initial learning rate $1\times 10^{-3}$. If the validation loss did not reach a new low for 5 consecutive epochs, the learning rate was halved. The loss function was defined as the scale-invariant signal-to-noise ratio (SI-SNR) [17] between the predicted and target speech signals.
@@
5.3 Comparison with Existing Methods
We conducted a systematic evaluation of Swift-Net on three datasets and compared it with various mainstream AVSS models. Results are shown in Table 1. To ensure fairness, all comparison methods were adapted to their causal versions following the causal strategy described in Section 3, and the open-source codes of each model were available [^2] https://github.com/JusperLee/Swift-Net. Moreover, all compared models were trained within the same training framework and evaluated using consistent performance metrics. Across all datasets, it was noteworthy that Swift-Net-6, as the fastest inference variant, achieved an SI-SNRi of 13.4 dB, outperforming existing SOTA causal methods such as causal AV-TF-GridNet. In addition, Swift-Net-6 reduced the parameter count by 11.8 times and the computational cost by 10.3 times compared with causal AV-TF-GridNet. Furthermore, our high-performance variant, Swift-Net-12, achieved over 1 dB improvement in performance relative to causal AV-TF-GridNet. These results collectively demonstrated that Swift-Net offered significant advantages for causal speech separation tasks.
In addition, we further conducted a comparative analysis of the separation performance of RTFSNet [31], AV-TF-GridNet [30], and Swift-Net in real-world multi-speaker scenarios. Specifically, we collected four distinct multi-speaker video samples from the YouTube platform. To facilitate intuitive comparison of the separated speech results, we developed an interactive demo website [^3] https://cslikai.cn/Swift-Net/ allowing users to listen and compare the outputs. Subjective listening evaluations indicated that, compared to RTFSNet and AV-TF-GridNet, Swift-Net produces speech signals that were clearer, more natural, effectively suppressing background interference.
@@ 5.4 Ablation Studies
Different Visual Feature Extraction Modules. We replaced the LightVid block visual‐feature encoder with those from ConvTasNet and RTFSNet while keeping every other component of the streaming AVSS pipeline identical in order to quantify its module‐level benefits. As shown in Table 2, this lean design requires only 67.27 K parameters and 3.39 M MACs – over 4× reduction relative to ConvTasNet and RTFSNet. Remarkably, despite its minimal complexity, LightVid block still delivers a slight performance edge, confirming that a structurally simple encoder can unite efficiency and effectiveness in real-time causal AVSS.
@@
Effect of Incorporating Power Spectrogram Features. We systematically evaluated the impact of introducing audio power spectrogram features on audio separation performance across three models: RTFSNet, AV-TF-GridNet, and Swift-Net. As shown in Table 6, in Swift-Net (with 0.53M parameters and an increase of less than 0.1G MACs), the inclusion of power spectrogram features raised the SI-SNRi from 13.1 to 13.3. Furthermore, introducing power spectrogram features resulted in negligible increases in model parameters and incurred minimal MACs overhead, making it a low-cost enhancement. We conjectured that power spectrogram features provided key speech-presence cues, thereby improving audio-visual fusion and separation performance.
Table 6: Results with (upper) and without (lower) power spectrogram modeling on the LRS2-2Mix dataset.
@@
RTFSNet
12.8
12.4
13.1
12.8
0.64
0.63
49.89
49.81
AV-TF-GridNet
13.0
12.7
13.4
13.1
6.26
6.25
213.96
213.88
Swift-Net-6
13.3
13.1
13.6
13.4
0.53
0.52
20.68
20.60